OpenStudy (anonymous):

Which function has the following characteristics? A vertical asymptote at x = 3 A horizontal asymptote at y = -1/2

OpenStudy (anonymous):

Looks like the (1/x) or (-1/x) graphs. To get the vertical asymptote, have the denominator such that when you equate the denominator to 0, x=3. i.e. x=3 ==> x-3 = 0. To get the horizontal asymptote, you'd have to shift the graph such that the new "y-axis" is -1/2. To do this, substitute y with y-1/2. 1) For 1/x: y-1/2 = 1 / (x-3) 2) For (-1/x): y-1/2 = -1 / (x-3)

OpenStudy (anonymous):

What? Im kinda confused.